OpenStudy (anonymous):

3/5 * 4/5 = A. 12/25 B. 3/4 C. 7/25 D. 7/10

OpenStudy (anonymous):

You will be multiplying 3/5 by 4/5 to get (3/5)(4/5) \[(\frac{ 3 }{ 5 })(\frac{ 4 }{ 5})\] Now you will simplify (3/5)(4/5) Multiply 3/5 and 4/5 to get \[\frac{ (3*4) }{ (5*5) }\] Multiply 3 by 4 to get 12. \[\frac{ 12 }{ 5*5 }\] Multiply 5 by 5 to get 25 \[\frac{ 12 }{ 25 }\] Exact Form: \[\frac{ 12 }{ 25 }\] Decimal Form; 0.48 @scotto11

OpenStudy (anonymous):

I hope this will help and make you understand about this question @scotto11